I have a few questions.
What would the monthly intrest rate be, for a $1000 dollar credit with a charge of $195 for a month?

Find two consecutive integers such that the sum of 2 times the first integer and 4 times second integer is 46.
Can you explain to me how this should be solved?
-3(10x-3)+10(3x-3)=x-7
convert 9/16 into a percent.

Answer by stanbon(75887)

What would the monthly interest rate be, for a $1000 dollar credit with a charge of $195 for a month?
195/1000= 0.195 = 19.5%
--------------------------------
Find two consecutive integers such that the sum of 2 times the first integer and 4 times second integer is 46.
Two consecutive integers are x and x+1:
EQUATION:
2x+4(x+1)=46
6x+4=46
6x=42
x=7 (1st integer)
x+1=8 (2nd integer)
-------------------------
Can you explain to me how this should be solved?
-3(10x-3)+10(3x-3)=x-7
-30x+9+30x-30=x-7
-21 = x-7
x=-14
---------------
convert 9/16 into a percent.
Percent means per-hundred, so write:
9/16 = x/100
x=900/16
x=56.25%
============
Cheers,
Stan H.